OpenStudy (anonymous):

Al saves pennies. Al agreed to give 3/5 of his pennies to Bev if she would give 3/5 of what she got from Al to Carl and if Carl in turn would give 3/5 of what he got from Bev to Dani. Bev, Carl, and Dani received 189 pennies. How many pennies did Al have initially?

OpenStudy (anonymous):

what's a penny? a coin? i don't know american coins. .

OpenStudy (anonymous):

a penny is one cent .01

OpenStudy (anonymous):

neways it makes no diff

OpenStudy (anonymous):

I gtg sorry i cant help out :(

OpenStudy (anonymous):

It is a coin.

OpenStudy (anonymous):

The answer was not correct. I'm sorry if you misunderstood, the question asks how many coins Al had to begin with.

OpenStudy (anonymous):

is it 189 each?

OpenStudy (anonymous):

does bec, carl and dani received 189 each?

OpenStudy (anonymous):

Click 0.JPG first then 00.JPG. This is an example only using different numbers. I am not understanding the example.

OpenStudy (anonymous):

If only Dani received 189 pennies, then Al started with 875. If all 3 of them combined received 189 pennies, Al had 315

Directrix (directrix):

875. (27/125) P = 189. P = 875

OpenStudy (anonymous):

let x be the amount Al started with x*3/5=amount Bev has x*3/5*3/5=amount Carl has x*3/5*3/5*3/5=amount Dani has so 189=x*3/5*3/5*3/5 x=875

OpenStudy (anonymous):

Thank you!! This was sucking the life out of me. ;D

OpenStudy (anonymous):

The wording of the problem is retarded. It makes it look like all 3 combined got 189.

OpenStudy (anonymous):

It totally is!

OpenStudy (anonymous):

3/5x + 3/5(3/5x) + 3/5[3/5(3/5x)]=x+189 is my representation correct??